OpenStudy (anonymous):

Given the function f(x) = 2(x + 4), find x if f(x) = 20.

OpenStudy (anonymous):

\[20 = 2(x+4) \rightarrow 20=2x+8\rightarrow12=2x \rightarrow x = 6\]

OpenStudy (anonymous):

It's not 48?

OpenStudy (anonymous):

If you plug in x =6 to the function you will see it comes out to 20.

OpenStudy (anonymous):

If you are trying to solve for f(20) then yes it's 48.

OpenStudy (anonymous):

Where did you get 6 from?

OpenStudy (anonymous):

look at the equation I showed you; all the steps are there

OpenStudy (anonymous):

\[20=2(x+4)→20=2x+8→12=2x→x=6\]

OpenStudy (anonymous):

\[f(x) = 2(x + 4)\]what the question is asking is to find what value of x would make the function equal 20, so x=? so that f(x)=20:\[20=2(x+4)\]\[20=2x+8\]\[12=2x\]\[6=x\]so when x=6, the function equals 20, so x=6 so that f(6)=20

OpenStudy (anonymous):

we can check that to be true:\[f(6)=2(6+4)\]\[f(6)=2(10)\]\[f(6)=20~\checkmark\]